# ROSE: Register Assisted General Time Series Forecasting with Decomposed Frequency Learning

### v1

@@ 4.2 Ablation studies
Replace Multi-Freq Masking Single Freq Masking 0.354 0.379 0.257 0.313 0.403 0.423 0.347 0.387
Multi Patch Masking 0.356 0.379 0.259 0.316 0.404 0.426 0.349 0.389
Patch Masking 0.378 0.4 0.261 0.319 0.408 0.432 0.375 0.407
w/o TS-Register 0.354 0.378 0.256 0.312 0.418 0.427 0.355 0.39
Predict Task 0.36 0.384 0.257 0.314 0.422 0.438 0.372 0.41
Reconstruct Task 0.387 0.403 0.269 0.327 0.412 0.428 0.361 0.399
From Scratch 0.370 0.393 0.261 0.318 0.470 0.480 0.401 0.425
4.3 Scalability
@@
Data size. When keeping the model size, we increase the size of the pre-training datasets from 178M to 886M. The results are shown in Figure 5(c) and (d). The performance of our model steadily improves with the increase in dataset size and achieves improvements of 7.4% and 4.8% respectively.
4.4 Model Analysis
Zero-shot and inference efficiency. By adopting the dual pre-train task, our model possesses strong generalization ability and can forecast without requiring any downstream data for fine-tuning.
Figure 6: Model performance in zero-shot setting comparison and inference efficiency comparison.
To validate ROSE’s zero-shot capabilities and assess its inference efficiency, we compared it with other models that have recently demonstrated similar abilities, namely Timer [2], MOIRAI [8], PatchTST [6], and GPT4TS [10]. Specific implementation details and results can be found in Appendix B. As shown in Figure 6, ROSE exhibits superior performance in zero-shot setting, with significantly fewer model parameters and inference time compared to other models. This underscores the robust generalization and scalability of ROSE.
Figure 7: Visualization of TS-Register.
@@ 4.4 Model Analysis
Sensitivity. The sensitivity analyses for the upper bound $a$ of the thresholds, the number of masked series $K_{\mathrm{f}}$, the number of register tokens $N_{\text{r}}$ and the size of codebook $H$ are presented in Appendix C.
5 Conclusion and future work
In this work, we propose ROSE, a novel general model, addressing the challenges of leveraging multi-domain datasets for enhancing downstream prediction task performance. ROSE utilizes decomposed frequency learning and TS-Register to capture generalized and domain-specific representations, enabling improved fine-tuning results, especially in data-scarce scenarios. Our experiments demonstrate ROSE’s superior performance over baselines with both full-data and few-data fine-tuning, as well as its impressive zero-shot capabilities. Future efforts will concentrate on expanding pre-training datasets and extending ROSE’s applicability across diverse time series analysis tasks, e.g., classification.
Appendix A IMPLEMENTATION DETAILS
A.1 Pre-training Datasets
We use multi-source datasets in pre-training which contain subsets of Monash [41], UEA [42] and UCR [43] time series regression datasets, as well as some other time series classical datasets [44, 45, 46, 47, 48]. The final list of all pre-training datasets is shown in Table 4. There is no overlap between the pre-training datasets and the target datasets. It is worth noting that the dataset weather in the pre-training dataset is a univariate dataset, which is different to the multivariate dataset weather in the target task. The pre-trained datasets can be categorized into 6 different domains according to their sources: Energy, Nature, Health, Transport, and Web. The sampling frequencies of the datasets show a remarkable diversity, ranging from millisecond samples to monthly samples, which reflects the diverse application scenarios and complexity of the real world. For all pre-training datasets, we split them into univariate sequences and train them in a channel-independent manner.
A.2 Evaluation Datasets
We use the following 8 multivariate time-series datasets for downstream fine-tuning and forecasting: ETT datasets contain 7 variates collected from two different electric transformers from July 2016 to July 2018. ETT datasets [^1] https://github.com/zhouhaoyi/ETDataset consist of four subsets, of which ETTh1/ETTh2 are recorded hourly and ETTm1/ETTm2 are recorded every 15 minutes. Traffic [^2] https://pems.dot.ca.gov/ contains road occupancy rates measured by 862 sensors on freeways in the San Francisco Bay Area from 2015 to 2016, recorded hourly. Weather [^3] https://www.bgc-jena.mpg.de/wetter/ collects 21 meteorological indicators, such as temperature and barometric pressure, for Germany in 2020, recorded every 10 minutes. Solar [^4] https://www.nrel.gov/grid/solar-power-data.html records solar power generation from 137 PV plants in 2006, every 10 minutes. Electricity [^5] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014 contains the electricity consumption of 321 customers from July 2016 to July 2019, recorded hourly. We split each evaluation dataset into train-validation-test sets and detailed statistics of evaluation datasets are shown in Table 3.
Table 3: The statistics of evaluation datasets
Dataset ETTm1 ETTm2 ETTh1 ETTh2 Traffic Weather Solar Electricity
Variables 7 7 7 7 862 21 137 321
Timestamps 69680 69680 17420 17420 17544 52696 52560 26304
Split Ratio 6:2:2 6:2:2 6:2:2 6:2:2 7:1:2 7:1:2 7:1:2 7:1:2
Table 4: List of pretraining datasets.
Domain Dataset Frequency Time Pionts Source
Energy Aus. Electricity Demand Half Hourly 1155264 Monash[41]
Wind 4 Seconds 7397147 Monash[41]
Wind Farms Minutely 172178060 Monash[41]
London Smart Meters Half Hourly 166527216 Monash[41]
Nature Phoneme - 2160640 UCR[43]
EigenWorms - 27947136 UEA[42]
PRSA Hourly 4628448 [44]
Temperature Rain Daily 23252200 Monash[41]
StarLightCurves - 9457664 UCR[43]
Worms 0.033 Seconds 232200 UCR[43]
Saugeen River Flow Daily 23741 Monash[41]
Sunspot Daily 73924 Monash[41]
Weather Daily 43032000 Monash[41]
KDD Cup 2018 Daily 2942364 Monash[41]
US Births Daily 7305 Monash[41]
Health MotorImagery 0.001 Seconds 72576000 UEA[42]
SelfRegulationSCP1 0.004 Seconds 3015936 UEA[42]
SelfRegulationSCP2 0.004 Seconds 3064320 UEA[42]
AtrialFibrillation 0.008 Seconds 38400 UEA[42]
PigArtPressure - 624000 UCR[43]
PIGCVP - 624000 UCR[43]
TDbrain 0.002 Seconds 79232703 [45]
Transport Pems03 5 Minute 9382464 [46]
Pems04 5 Minute 5216544 [46]
Pems07 5 Minute 24921792 [46]
Pems08 5 Minute 3035520 [46]
Pems-bay 5 Minute 16937700 [46]
Pedestrian_Counts Hourly 3132346 Monash[41]
Web Web Traffic Daily 116485589 Monash[41]
Economic FRED_MD Monthly 77896 [47]
Bitcoin Daily 75364 Monash[41]
NN5 Daily 87801 [48]
A.3 Setting
We implemented ROSE in PyTorch [49] and all the experiments were conducted on 8 NVIDIA A800 80GB GPU. We used ADAM [50] with an initial learning rate of $5\times 10^{-4}$ and implemented learning rate decay using the StepLR method to implement learning rate decaying pre-training. By default, ROSE contains 3 encoder layers and 3 decoder layers with head number $H=16$ and the dimension of latent space $D=256$.
Pre-training. We use $N_{\mathrm{r}}=3$ as the number of register tokens and $P=8$ as the path tokens. We set the input length to 512 for the supervised prediction task with target lengths of 96, 192, 336, and 720. We also set the input length to 512 for self-supervised multi-frequency mask-reconstruction task, where mask number $K_{\mathbf{f}}=4$. The batch size is set to 8192 in pre-training.
Fine-tuning. We fix the lookback window to 512, and perform predictions with target lengths of 96, 192, 336, and 720, respectively. The number of register tokens $N_{\mathrm{r}}$ and patch tokens $P$ is the same as in pre-training, and the parameter $k=3$ in topk is set when selection vectors are performed in the register codebook. The batch size is set to 64 in fine-tuning.
A.4 Baselines
We select the state-of-the-art methods as our baselines to compare the performance of different models with full-data fine-tuning and 10% few-shot setting, including Transformer-based methods: Timer [2], iTransformer [20], and PatchTST [6]; CNN-based backbones method: TimesNet [18]; MLP-based methods: FITS [38], TIDE [39], and DLinear [40]. It is worth noting that Timer is a recent general time series model with a GPT-style architecture. Since the size of the look-back window can affect the performance of different models, we choose the look-back window size in $\{96,336,512,720\}$ for each baseline that achieves its best results for fair comparisons. The specific code base for these models is listed in Table 5:
Table 5: Code repositories for baselines
Baselines code repositories
Timer https://github.com/thuml/Large-Time-Series-Model
iTransformer https://github.com/thuml/iTransformer
PatchTST https://github.com/yuqinie98/PatchTST
FITS https://anonymous.4open.science/r/FITS/README.md
TimesNet https://github.com/thuml/TimesNet
TIDE https://github.com/google-research/google-research/tree/master/tide
Dlinear https://github.com/cure-lab/LTSF-Linear
Appendix B Zero shot and inference efficiency
To explore the inference efficiency and zero-shot capability of ROSE, we selected recent models with zero-shot capability, including Timer [2] and MOIRAI [8], as well as models with the potential for zero-shot learning, including PatchTST [6] and GPT4TS [10], for comparison on the ETT dataset. And we evaluate their inference efficiency by recording inference time per batch on the ETTh2 dataset with a batch size of 32.
Pre-train: We select the models with the largest parameter weights for Timer and MOIRAI as baselines. GPT4TS and PatchTST were pre-trained on our pre-training dataset with prediction tasks to acquire zero-shot capability. The specific settings are shown in the table 6.
Detailed results: For fare comparison, we set the input length as 96, 336, 512, and 720 for all zero-shot baseline models. We do not consider longer input length, as many real-world scenarios could offer very limited samples, and thus the zero-shot setting with only a few samples is crucial for them. As the baselines do not contain results with this setting in their original papers, we rerun them and report their best results in Table 11. ROSE consistently surpasses the baselines despite having significantly fewer parameters and inference time.
Table 6: Pre-train setting and inference efficiency
Model Parameters Pre-train data size Inference time per batch code repositories
ROSE 6.4M 0.89B 4.3 ms -
MOIRAI_large 311M 27B 90.3ms https://github.com/redoules/moirai
GPT4TS 81.9M 0.89B 45.6 ms https://github.com/DAMO-DI-ML/NeurIPS2023-One-Fits-All
Timer 67.4M 1B 68 ms https://github.com/thuml/Large-Time-Series-Model
PatchTST 5.9M 0.89B 4.2 ms https://github.com/PatchTST/PatchTST
Appendix C Sensitivity
Number of masked series. As described in Section 3.2, we propose decomposed frequency learning, which employs multiple thresholds to randomly mask high and low frequencies in the frequency domain, thereby decomposing the original time series into multiple frequency components. This allows the model to understand the time series from multiple frequency perspectives. In this experiment, we study the influence of the number of masked series $K_{m}athrm{f}$ on downstream performance. We train ROSE with 1, 2, 3, 4, 5, or 6 mask series. We report the results of this analysis in Figure 8(a). We find that as the number of masked sequences increases, the downstream performance gradually improves. This is because the model can better understand the time series from the decomposed frequency components, which enhances the model’s generalization ability. However, more masked series do not bring better downstream performance. This could be due to an excessive number of masked sequences leading to information redundancy. In all our experiments, we kept 4 mask series.
@@ Appendix C Sensitivity
Figure 8: (a): Analysis of the number of masked series. (b): Analysis of the number of register tokens
@@
Register codebook size. Figure 9(b) illustrates the relationship between register codebook size and model performance. The codebook size determines the upper limit of domain-specific information that the register can store. We can observe that there is a significant improvement in the model effect when the codebook size is increased from 32 to 128. When the codebook size exceeds 128, the improvement of the model effect with the increase of codebook size is no longer obvious. Therefore, we believe that 128 is an appropriate codebook size for the current pre-training datasets.
Figure 9: (a): Analysis of the threshold upper bound. (b): Analysis of the size of codebook
Appendix D Model generality
We evaluate the effectiveness of our proposed Multiple Frequency Mask (MFM) on Transformer-based models and CNN-based models, whose results are shown in Table 7. It is notable that MFM consistently improves these forecasting models. Specifically, it achieves average improvements of 6.3%, 3.7%, 1.5% in Autoformer [19], TimesNet [18], and PatchTST [6], respectively. This indicates that MFM can be widely utilized across various time series forecasting models to learn generalized time series representations and improve prediction accuracy.
Table 7: Performance Multiple Frequency Masking.

### v2

@@ 4.4 Ablation studies
w/o TS-Register 0.354 0.378 0.256 0.312 0.418 0.427 0.355 0.390
Prediction Task 0.360 0.384 0.257 0.314 0.422 0.438 0.372 0.410
Reconstruction Task 0.387 0.403 0.269 0.327 0.412 0.428 0.361 0.399
From scratch 0.371 0.391 0.261 0.318 0.470 0.480 0.400 0.425
Table 4: Ablations on decomposed frequency learning, where we replace Multi-freq masking with other masking methods. The average results of all predicted lengths are listed here.
Design ETTm1 ETTm2 ETTh1 ETTh2
MSE MAE MSE MAE MSE MAE MSE MAE
ROSE 0.349 0.372 0.250 0.308 0.397 0.419 0.335 0.380
Replace Multi-Frequency Masking Random Freq Masking 0.381 0.397 0.261 0.324 0.410 0.427 0.374 0.405
Multi-Patch Masking 0.356 0.379 0.259 0.316 0.404 0.426 0.349 0.389
Patch Masking 0.378 0.400 0.261 0.319 0.408 0.432 0.375 0.407
From scratch 0.371 0.391 0.261 0.318 0.470 0.480 0.400 0.425
5 Conclusion and future work
In this work, we propose ROSE, a novel general model, addressing the challenges of leveraging multi-domain datasets for enhancing downstream prediction task performance. ROSE utilizes decomposed frequency learning and TS-Register to capture generalized and domain-specific representations, enabling improved fine-tuning results, especially in data-scarce scenarios. Our experiments demonstrate ROSE’s superior performance over baselines with both full-data and few-data fine-tuning, as well as its impressive zero-shot capabilities. Future efforts will concentrate on expanding pre-training datasets and extending ROSE’s applicability across diverse time series analysis tasks, e.g., classification. We provide our code at https://anonymous.4open.science/r/ROSE-A235.
6 Reproducibility
Our work meets reproducibility requirements. Specifically, you can download our evaluation datasets in a standardized format from the public link (Wang et al., 2024a): https://drive.google.com/drive/folders/13Cg1KYOlzM5C7K8gK8NfC-F3EYxkM3D2, and we provide our code at https://anonymous.4open.science/r/ROSE-A235.
Appendix A Appendix
A.1 IMPLEMENTATION DETAILS
A.1.1 Pre-training Datasets
We use multi-source datasets in pre-training which contain subsets of Monash [Godahewa et al., 2021], UEA [Bagnall et al., 2018] and UCR [Dau et al., 2019] time series datasets, as well as some other time series classical datasets [Zhang et al., 2017, Wang et al., 2024b, Liu et al., 2022b, McCracken and Ng, 2016, Taieb et al., 2012]. The final list of all pre-training datasets is shown in Table 6. There is no overlap between the pre-training datasets and the target datasets. It is worth noting that the dataset weather in the pre-training dataset is a univariate dataset, which is different to the multivariate dataset weather in the target task. The pre-trained datasets can be categorized into 6 different domains according to their sources: Energy, Nature, Health, Transport, and Web. The sampling frequencies of the datasets show a remarkable diversity, ranging from millisecond samples to monthly samples, which reflects the diverse application scenarios and complexity of the real world. For all pre-training datasets, we split them into univariate sequences and train them in a channel-independent manner.
A.1.2 Evaluation Datasets
We use the following 7 multivariate time-series datasets for downstream fine-tuning and forecasting: ETT datasets [^1] https://github.com/zhouhaoyi/ETDataset contain 7 variates collected from two different electric transformers from July 2016 to July 2018. It consists of four subsets, of which ETTh1/ETTh2 are recorded hourly and ETTm1/ETTm2 are recorded every 15 minutes. Traffic [^2] https://pems.dot.ca.gov/ contains road occupancy rates measured by 862 sensors on freeways in the San Francisco Bay Area from 2015 to 2016, recorded hourly. Weather [^3] https://www.bgc-jena.mpg.de/wetter/ collects 21 meteorological indicators, such as temperature and barometric pressure, for Germany in 2020, recorded every 10 minutes. Electricity [^4] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014 contains the electricity consumption of 321 customers from July 2016 to July 2019, recorded hourly. We split each evaluation dataset into train-validation-test sets and detailed statistics of evaluation datasets are shown in Table 5.
Table 5: The statistics of evaluation datasets.
Dataset ETTm1 ETTm2 ETTh1 ETTh2 Traffic Weather Electricity
Variables 7 7 7 7 862 21 321
Timestamps 69680 69680 17420 17420 17544 52696 26304
Split Ratio 6:2:2 6:2:2 6:2:2 6:2:2 7:1:2 7:1:2 7:1:2
Table 6: List of pretraining datasets.
Domain Dataset Frequency Time Pionts Source
Energy Aus. Electricity Demand Half Hourly 1155264 Monash[Godahewa et al., 2021]
Wind 4 Seconds 7397147 Monash[Godahewa et al., 2021]
Wind Farms Minutely 172178060 Monash[Godahewa et al., 2021]
Solar 10 Minutes 7200720 Monash[Godahewa et al., 2021]
Solar Power 4 Seconds 7397222 Monash[Godahewa et al., 2021]
London Smart Meters Half Hourly 166527216 Monash[Godahewa et al., 2021]
Nature Phoneme - 2160640 UCRDau et al. [2019]
EigenWorms - 27947136 UEA[Bagnall et al., 2018]
PRSA Hourly 4628448 [Zhang et al., 2017]
Temperature Rain Daily 23252200 Monash[Godahewa et al., 2021]
StarLightCurves - 9457664 UCR[Dau et al., 2019]
Worms 0.033 Seconds 232200 UCR[Dau et al., 2019]
Saugeen River Flow Daily 23741 Monash[Godahewa et al., 2021]
Sunspot Daily 73924 Monash[Godahewa et al., 2021]
Weather Daily 43032000 Monash[Godahewa et al., 2021]
KDD Cup 2018 Daily 2942364 MonashGodahewa et al. [2021]
US Births Daily 7305 Monash[Godahewa et al., 2021]
Health MotorImagery 0.001 Seconds 72576000 UEA[Bagnall et al., 2018]
SelfRegulationSCP1 0.004 Seconds 3015936 UEA[Bagnall et al., 2018]
SelfRegulationSCP2 0.004 Seconds 3064320 UEA[Bagnall et al., 2018]
AtrialFibrillation 0.008 Seconds 38400 UEA[Bagnall et al., 2018]
PigArtPressure - 624000 UCR[Dau et al., 2019]
PIGCVP - 624000 UCR[Dau et al., 2019]
TDbrain 0.002 Seconds 79232703 [Wang et al., 2024b]
Transport Pems03 5 Minute 9382464 [Liu et al., 2022b]
Pems04 5 Minute 5216544 [Liu et al., 2022b]
Pems07 5 Minute 24921792 [Liu et al., 2022b]
Pems08 5 Minute 3035520 [Liu et al., 2022b]
Pems-bay 5 Minute 16937700 [Liu et al., 2022b]
Pedestrian_Counts Hourly 3132346 Monash[Godahewa et al., 2021]
Web Web Traffic Daily 116485589 Monash[Godahewa et al., 2021]
Economic FRED_MD Monthly 77896 [McCracken and Ng, 2016]
Bitcoin Daily 75364 Monash[Godahewa et al., 2021]
NN5 Daily 87801 [Taieb et al., 2012]
A.1.3 Setting
We implemented ROSE in PyTorch [Paszke et al., 2019] and all the experiments were conducted on 8 NVIDIA A800 80GB GPU. We used ADAM [Kingma and Ba, 2014] with an initial learning rate of $5\times 10^{-4}$ and implemented learning rate decay using the StepLR method to implement learning rate decaying pre-training. By default, ROSE contains 3 encoder layers and 3 decoder layers with head number of 16 and the dimension of latent space $D=256$. The patch size for patching is set to 64.
Pre-training. We use $N_{\mathrm{r}}=3$ as the number of register tokens and $P=8$ as the path tokens. We set the input length to 512 for the supervised prediction task with target lengths of 96, 192, 336, and 720. We also set the input length to 512 and mask number $K_{\mathbf{f}}=4$. The batch size is set to 8192 in pre-training.
Fine-tuning. We fix the lookback window to 512, and perform predictions with target lengths of 96, 192, 336, and 720, respectively. The number of register tokens $N_{\mathrm{r}}$ and patch tokens $P$ is the same as in pre-training, and the parameter $k=3$ in TopK is set when selection vectors are performed in the register.
A.1.4 Baselines
We select the state-of-the-art models as our baselines in full-shot and few-shot setting, including four specific models: iTransformer [Liu et al., 2023], PatchTST [Nie et al., 2022], TimesNet [Wu et al., 2022], and DLinear [Zeng et al., 2023], and two LLM-based models: GPT4TS [Zhou et al., 2024] and $\textbf{S}^{2}$-LLM [Pan et al., 2024]. In addition, we selected five foundation models for comparison in zero-shot setting, including Timer [Liu et al., 2024], MOIRAI [Woo et al., 2024], Chronos [Ansari et al., 2024], TimesFM [Das et al., 2023b] and Moment [Goswami et al., 2024]. The specific code base for these models is listed in Table 7:
Table 7: Code repositories for baselines.
Model Types Models Code Repositories
Small Model iTransformer https://github.com/thuml/iTransformer
PatchTST https://github.com/yuqinie98/PatchTST
TimesNet https://github.com/thuml/TimesNet
Dlinear https://github.com/cure-lab/LTSF-Linear
Foundation Model Timer https://github.com/thuml/Large-Time-Series-Model
MOIRAI https://github.com/redoules/moirai
Chronos https://github.com/amazon-science/chronos-forecasting
TimesFM https://github.com/google-research/timesfm/
Moment https://anonymous.4open.science/r/BETT-773F/README.md
LLM-based Model GPT4TS https://github.com/DAMO-DI-ML/NeurIPS2023-One-Fits-All
S2P-LLM https://github.com/panzijie825/S2IP-LLM
A.2 Efficiency analysis
As an important aspect of foundation models, inference efficiency is crucial. Therefore, we evaluate the testing time of ROSE and five foundation models in the ETTh1 and ETTh2 dataset in zero-shot setting. Similarly, we evaluate the time of the entire process of training, validation, and testing for four specific models in the same datasets in full-shot setting. The above experiments all set the batch size to 32. The specific results are shown in Table 8 and Figure 5. We observe that ROSE maintains its advantage in zero-shot performance while also being significantly faster compared to the baselines, even being approximately ten times faster than the second-fastest foundation model, Timer. This raises our reflection on whether time-series foundation models require extremely large parameter sizes and whether existing time-series foundation models have validated their architectures’ scaling laws on time-series data.
Table 8: Efficiency analysis.
Model Parameters Pre-train datasize Averaged time
ROSE 7.4M 0.89B 0.652s
MOIRAI 311M 27B 7.920s
Timer 67.4M 1B 5.989s
Chronos 46M 84B 176s
TimesFM 200M 100B 167.2s
Moment 385M 1.13B 88s
Itransformer 3.8M - 34.18s
PatchTST 3.2M - 35.47s
TimesNet 1.8M - 146s
Dlinear 0.018M - 24.06s
A.3 Calculation of cosine similarity
In Figure 6, we visualize the cosine similarity of register vector selections. For each sample in a dataset, during the inference process, $k$ vectors are selected from the register based on the Top-K strategy. We iterate through all samples in the dataset and count the number of times each vector is selected, which allows us to obtain a record vector of length equivalent to the size of the register for each dataset. The $i_{th}$ position in the record vector represents the number of times the $i_{th}$ vector in the register has been selected by samples in the dataset. For a pair of datasets, we can obtain a unique record vector for each dataset, and then we are able to calculate the cosine similarity of two vectors.
A.4 Scalability
@@
Data size. When keeping the model size, we increase the size of the pre-training datasets from 178M to 887M. The results are shown in Figure 7(c) and (d). The performance of our model steadily improves with the increase in dataset size and achieves improvements of 7.4% and 4.8% respectively.
A.5 Sensitivity
@@
We perform the sensitivity analyses for the upper bound $a$ of the thresholds, the number of masked series $K_{\mathrm{f}}$, the number of register tokens $N_{\text{r}}$, the size of register $H$ and the number of selections $k$ in Top-K strategy. All the sensitivity experiments present the average results on the four ETT datasets: ETTh1, ETTh2, ETTm1 and ETTm2 under 10% few-shot setting.
Number of masked series. As described in Section 3.2, we propose decomposed frequency learning, which employs multiple thresholds to randomly mask high and low frequencies in the frequency domain, thereby decomposing the original time series into multiple frequency components. This allows the model to understand the time series from multiple frequency perspectives. In this experiment, we study the influence of the number of masked series $K_{\mathrm{f}}$ on downstream performance. We train ROSE with 1, 2, 3, 4, 5, or 6 mask series. We report the results of this analysis in Figure 8(a). We find that as the number of masked sequences increases, the downstream performance gradually improves. This is because the model can better understand the time series from the decomposed frequency components, which enhances the model’s generalization ability. However, more masked series do not bring better downstream performance. This could be due to an excessive number of masked sequences leading to information redundancy. In all our experiments, we keep 4 mask series.
@@
Figure 8: (a): Analysis of the number of masked series. (b): Analysis of the number of register tokens.
@@
Register size. Figure 9(b) illustrates the relationship between register size and model performance. The register size determines the upper limit of domain-specific information that the register can store. We can observe that there is a significant improvement in the model effect when the register size is increased from 32 to 128. When the register size exceeds 128, the improvement of the model effect with the increase of register size is no longer obvious. Therefore, we believe that 128 is an appropriate register size for the current pre-training datasets.
Number of selections in Top-K strategy. Figure 9(c) illustrates the relationship between the number of selections $k$ in Top-K strategy and model performance when we use the register to realize adaptive transfer of domain-specific information in downstream tasks. It can be seen that the model effect performance peaks at 3 tokens at $k$ = 3, which has some advantages over selecting once ($k$ = 1), indicating that the TopK strategy can compensate for the problem of incomplete matching of upstream and downstream domains to some extent. However, too large $k$ will also introduce redundant information and limit the accuracy of domain-specific information transfer.
Figure 9: (a): Analysis of the threshold upper bound. (b): Analysis of the size of register. (c): Analysis of the number of selections in Top-K strategy.
A.6 Short-term forecasting
We also try to apply ROSE to short-term forecasting on the M4 [Makridakis, 2018] dataset, which contains the yearly, quarterly and monthly collected univariate marketing data. We follow TimesNet’s [Wu et al., 2022] setting and metrics (SMAPE, MASE and OWA) for testing. As shown in Table 9, ROSE also exhibits competitive performance on the M4 dataset compared to the baselines.
Table 9: Full results on Short-term forecasting.
Metric ROSE iTransformer PatchTST Timesnet Dlinear GPT4TS S2IP-LLM
SMAPE MASE OWA SMAPE MASE OWA SMAPE MASE OWA SMAPE MASE OWA SMAPE MASE OWA SMAPE MASE OWA SMAPE MASE OWA
Yearly 13.302 3.014 0.833 13.238 2.952 0.823 16.766 4.331 1.018 13.387 2.996 0.786 16.965 4.283 1.058 13.531 3.015 0.793 13.413 3.024 0.792
Quarterly 9.998 1.165 0.885 10.001 1.278 0.949 12.132 1.513 0.966 10.100 1.182 0.890 12.145 1.520 1.106 10.100 1.194 0.898 10.352 1.228 0.922
Monthly 12.650 0.915 0.866 13.399 1.031 0.949 13.428 0.997 0.948 12.670 0.933 0.933 13.514 1.037 0.956 12.894 0.956 0.897 12.995 0.970 0.910
Others 4.668 3.126 1.020 6.558 4.511 1.401 6.667 4.834 1.417 4.891 3.302 1.035 6.709 4.953 1.487 4.940 3.228 1.029 4.805 3.247 1.071
A.7 Model generality
We evaluate the effectiveness of our proposed multi-frequency masking on Transformer-based models and CNN-based models, whose results are shown in Table 10. It is notable that multi-frequency masking consistently improves these forecasting models. Specifically, it achieves average improvements of 6.3%, 3.7%, 1.5% in Autoformer [Wu et al., 2021], TimesNet [Wu et al., 2022], and PatchTST [Nie et al., 2022], respectively. This indicates that multi-frequency Masking can be widely utilized across various time series forecasting models to learn generalized time series representations and improve prediction accuracy.
Table 10: Performance of multi-frequency masking.